# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

### 01 Feb 2024 v1

@@ Abstract
Figure 1: Our neuro-symbolic framework enables a seamless transition between symbolic and differentiable programming, each with distinct dynamics and strengths. Differentiable programming provides access to foundational and specialist models. Classical programming, on the other hand, shifts between abstraction and implementation, focusing on high-level concepts before delving into the details of implementation.
@@ 1 Introduction
The recent surge in generative AI, particularly involving large language models (LLMs), has demonstrated their wide-ranging applicability across various domains (Badita, 2022; Degrave, 2022). These models have enhanced the functionality of tools for search-based interactions (Microsoft, 2023; YouWrite, 2022; Writesonic, 2022), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more. Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few. Therefore, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023). Yet, these applications merely scratch the surface of the transformation that language-based interactions are expected to bring to human-computer interactions in both the near and distant future.
@@
In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate compositional, multi-modal, and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023). SymbolicAI augments the generative process of LLMs with functional zero- and few-shot learning operations and enables the creation of versatile applications through in-context learning (Wei et al., 2022a). These operations guide the generative process and facilitate a modular design with a wide range of existing solvers, including formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval. It exposes these solvers as building blocks for constructing compositional functions as computational graphs, and facilitates the development of an extensible toolkit that bridges classical and differentiable programming paradigms, aiming to create domain-invariant problem solvers. In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Deniz et al., 2019; Hu et al., 2022; Menenti et al., 2011; Scott et al., 2016; Macsweeney, 2002; Fedorenko et al., 2010; Regev et al., 2013), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016). We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022, 2023), that defines a stepping stone towards broad AI systems (see Section B).
@@ 3 Problem Definition
Conventional approaches employing foundation models for inference, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted in-context learning prompt instructions. This restricted scope limits the utilization to single modalities, lacks refinement or verification, and exhibits limited tool proficiency. We posit that the integration of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective. This approach has several advantages. Firstly, it facilitates the integration of pre-existing engineered solutions (e.g. various classical algorithms), offloading computational complexity and bridging various modalities. Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making. Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems. Central to our solution is a method to define and measure the orchestration of interactions between symbolic and sub-symbolic systems, and the level at which instructions are formulated for effective control and task execution.
@@ 4 Design Principles
In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines.
@@
Recently, several in-context learning methodologies evolved to enable tool usage through LLMs (Schick et al., 2023), or refine the generative outcome of LLMs (Yang et al., 2023). This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Singhal et al., 2023; Wei et al., 2022b). CoT prompting breaks down complex tasks into simpler, sequential steps, and helps with interpreting LLM’s output. Self-generated CoT, where models are encouraged to generate their own reasoning chains based on training examples, surpasses even expertly crafted CoT (Fernando et al., 2023). This observation echoes other reports that GPT-4 has an emergent self-improving capability through introspection, such as self-verification (Weng et al., 2023) or self-consistency (Wang et al., 2023b). Tree of Thoughts (ToT) enables LLMs to solve complex problems by exploring multiple reasoning paths through a search tree of coherent text units, demonstrating significant problem-solving enhancements in tasks requiring strategic planning and search (Yao et al., 2023a). Ensemble techniques further enhance the robustness and accuracy of model predictions by combining several strategies to establish a consensus (Nori et al., 2023). Conceptually, we designed our framework to enable all these techniques and combine them into dedicated components and sub-processes. Figure 2 outlines a typical expression evaluation in our NeSy pipeline.
Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through NeSy engines. The expression showcases the left shift operator $\ll$ and how the information of the symbol B is included in the symbol AC. The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below that represent task-specific few-shot examples.
@@
In-context learning enabled LLMs to become versatile task solvers by interpolating within the training distribution, to the extent that even potentially unseen tasks are addressable (Brown et al., 2020). We attribute this to associations formed within the input space and the capacity of Transformer architectures for defining domain-invariant feature sub-spaces. This phenomenon has striking parallels with advancements in few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks, that effectively acts to regularize the learning process, representing features that are invariant across different learning tasks. Furthermore, SubGD reflects the potential of in-context learning when combined with task-specific fine-tuning by showing that fine-tuning within a learned subspace significantly outperforms traditional fine-tuning methods. We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when use alongside with task-specific fine-tuning and solvers. To develop learning and reasoning systems capable of general problem-solving, we adopt a hybrid methodology. This approach leverages the in-context generalization capability of LLMs, constructing symbolic associations that aim to preserve and propagate situational context, and validating solutions with established solvers.
@@ 5 Framework
Figure 3: The illustrated pipeline utilizes GPT-4 as a NeSy engine and maps all user input and object instances to prompts. Description from left to right; yellow marks user data, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder. The user arguments and Symbol instance values are evaluated and transformed according to the pre- and post-processing steps. A custom method is constructed by using a reference implementation of decorators and passes its user arguments and instance attributes to the respective engine. A pre-processor manipulates the representation of user arguments and Symbol instance attributes before forwarding them to the engine. For instance, in case the engine is an LLM (i.e. NeSy engine), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine. Other attributes, such as examples or templates, can help pre-define structures for the desired LLM output format. After the result is computed by the engine, it is passed to the post-processors and validated against the constraints.
@@ 6 Evaluation
Figure 5: We evaluate GPT-3.5, GPT4, Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs). We denote the similarity score such that higher is better. The similarity score is measured according to a reference baseline and normalized to exclude noise.
@@ 7.2 Future Work
The goal for Algorithm 1 is to be utilized by an advanced learning agent. This agent, employing reinforcement learning methodologies (Ouyang et al., 2022; Li et al., 2023; Rafailov et al., 2023), could leverage our evaluation measure in existing benchmarks (Milani et al., 2020; Swazinna et al., 2022; Schweighofer et al., 2022) as a means to obtain reward signals to addresses a central problem in reinforcement learning, namely credit assignment (Sutton, 1984; Arjona-Medina et al., 2019; Holzleitner et al., 2020; Patil et al., 2020; Widrich et al., 2021; Dinu et al., 2022). Over time, it aims to develop the ability to autonomously generate its own plans, efficiently schedule tasks and subtasks, and carefully select the most suitable tools for each task. Our protocol lays the groundwork for this agent to learn and expand its base set of capabilities (Amaro et al., 2023), moving towards more sophisticated, self-referential orchestration of multi-step tasks. We’ve already noticed that research is shifting towards this type of methodology (Yuan et al., 2024). Furthermore, in Section 6 we’ve only considered a sequential scheduler. However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution.
@@ 7.3 Broader Impact
Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure. Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs. This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time. We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023). We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing artificial general intelligent agents.
@@ Acknowledgement
We are also grateful to Tim Scarfe, a friend whose community has been a hub for exhilarating discussions. His online presence and engagement have enriched the AI research landscape and broadened our perspectives.
@@ Appendix B Broad AI and Neuro-Symbolic Systems
Figure 6: Integration of symbolic and sub-symbolic AI in broad AI Systems, showcasing central concepts of broad AI systems, including knowledge, adaptability, abstraction, interaction, interpretability, efficiency, robustness, and reasoning (Hochreiter, 2022).
Our primary objective is to combine the strengths of symbolic and sub-symbolic approaches to overcome individual limitations. Symbolic AI is characterized by its emphasis on knowledge representation, the ability to abstract and formulate mathematical concepts, and the capacity for interactions with users or other systems in a human-understandable manner. These attributes ensure that we develop reasoning-based, interpretable AI systems with innate robustness and trustworthiness (Winter et al., 2021).
Both methodologies form the basis for developing new cognitive architectures, as illustrated in Figure 7. This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.

### 05 Feb 2024 v2

@@ Abstract
Figure 1: Our neuro-symbolic framework enables a seamless transition between classical and differentiable programming, each with distinct dynamics and strengths. Differentiable programming provides access to foundational and specialist models. Classical programming, on the other hand, shifts between abstraction and implementation, focusing on high-level concepts before delving into the details of implementation.
@@ 1 Introduction
The recent surge in generative AI, particularly involving large language models (LLMs), has demonstrated their wide-ranging applicability across various domains (Badita, 2022; Degrave, 2022). These models have enhanced the functionality of tools for search-based interactions (YouWrite, 2022; Writesonic, 2022; Microsoft, 2023), program synthesis (Jain et al., 2021; Romera-Paredes et al., 2023; Key et al., 2023), chat-based interactions (ReplikaAI, 2016; OpenAI, 2022; Google, 2023), and many more. Moreover, language-based approaches have facilitated connections between different modalities, enabling text-to-image (Ramesh et al., 2021; Saharia et al., 2022), text-to-video (Singer et al., 2022), text-to-3D (Poole et al., 2022), text-to-audio (Oord et al., 2016; Wang et al., 2017), and text-to-code (Wang et al., 2021b; Lu et al., 2021; Li et al., 2022b) transformations, to name a few. Therefore, by training on vast quantities of unlabelled textual data, LLMs have been shown to not only store factual knowledge (Petroni et al., 2019; Kassner et al., 2020) and approximate users’ intentions to some extent (Andreas, 2022), but also to unlock deep specialist capabilities through innovative prompting techniques (Nori et al., 2023). Yet, these applications merely scratch the surface of the transformation that language-based interactions are expected to bring to human-computer interactions in both the near and distant future.
@@
In light of these considerations, we introduce SymbolicAI, a compositional neuro-symbolic (NeSy) framework able to represent and manipulate multi-modal and self-referential structures (Schmidhuber, 2007; Fernando et al., 2023). SymbolicAI augments the generative process of LLMs with functional zero- and few-shot learning operations and enables the creation of versatile applications through in-context learning (Wei et al., 2022a). These operations guide the generative process and facilitate a modular design with a wide range of existing solvers, including formal language engines for mathematical expression evaluation, theorem provers, knowledge bases, and search engines for information retrieval. It exposes these solvers as building blocks for constructing compositional functions as computational graphs, and facilitates the development of an extensible toolkit that bridges classical and differentiable programming paradigms, aiming to create domain-invariant problem solvers. In designing the architecture of SymbolicAI, we drew inspiration from the body of evidence that suggests the human brain possesses a selective language processing module (Macsweeney, 2002; Fedorenko et al., 2010; Menenti et al., 2011; Regev et al., 2013; Scott et al., 2016; Deniz et al., 2019; Hu et al., 2022), prior research on cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Laird, 2022), and the significance of language on the structure of semantic maps in the human brain (Huth et al., 2016). We consider language as a central processing module, distinct from other cognitive processes such as reasoning or memory (Paischer et al., 2022, 2023), that defines a stepping stone towards broad AI systems (see Section B).
@@ 3 Problem Definition
Conventional approaches employing foundation models for inference, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted in-context learning prompt instructions. This restricted scope limits the utilization to single modalities, lacks refinement or verification, and exhibits limited tool proficiency. We posit that the integration of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective. This approach has several advantages. Firstly, it facilitates the integration of pre-existing engineered solutions (e.g. various classical algorithms), offloading computational complexity and bridging various modalities. Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making (e.g. selecting the respective tool based on in-context classification). Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems. Central to our solution is a method to define and measure the orchestration of interactions between symbolic and sub-symbolic systems, and the level at which instructions are formulated for effective control and task execution.
@@ 4 Design Principles
In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines. For further technical details, see Section 5.
@@
Recently, several in-context learning methodologies evolved to enable tool usage through LLMs (Schick et al., 2023), or refine the generative outcome of LLMs (Yang et al., 2023). This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Wei et al., 2022b; Singhal et al., 2023). CoT prompting breaks down complex tasks into simpler, sequential steps, and helps with interpreting LLM’s output. Self-generated CoT, where models are encouraged to generate their own reasoning chains based on training examples, surpasses even expertly crafted CoT (Fernando et al., 2023). This observation echoes other reports that GPT-4 has an emergent self-improving capability through introspection, such as self-verification (Weng et al., 2023) or self-consistency (Wang et al., 2023b). Tree of Thoughts (ToT) enables LLMs to solve complex problems by exploring multiple reasoning paths through a search tree of coherent text units, demonstrating significant problem-solving enhancements in tasks requiring strategic planning and search (Yao et al., 2023a). Ensemble techniques further enhance the robustness and accuracy of model predictions by combining several strategies to establish a consensus (Nori et al., 2023). Conceptually, we designed our framework to enable all these techniques and combine them into dedicated components and sub-processes. Figure 2 outlines a typical expression evaluation in our NeSy pipeline.
Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through NeSy engines. The expression showcases the left shift operator $\ll$ and how the information of the symbol B is included in the symbol AC. The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below represent task-specific few-shot examples.
@@
In-context learning enabled LLMs to become versatile task solvers by interpolating within the training distribution, to the extent that even potentially unseen tasks are addressable (Brown et al., 2020). We attribute this to associations formed within the input space and the capacity of Transformer architectures for defining domain-invariant feature sub-spaces. This phenomenon has striking parallels with advancements in few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks, that effectively acts to regularize the learning process, representing features that are invariant across different learning tasks. Furthermore, SubGD reflects the potential of in-context learning when combined with task-specific fine-tuning by showing that fine-tuning within a learned subspace significantly outperforms traditional fine-tuning methods. We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when used alongside with task-specific fine-tuning and solvers. To develop learning and reasoning systems capable of general problem-solving, we adopt a hybrid methodology. This approach leverages the in-context generalization capability of LLMs, constructing symbolic associations that aim to preserve and propagate situational context, and validating solutions with established solvers.
@@ 5 Framework
Figure 3: We illustrate our NeSy engine pipeline and how we map user input and object instance attributes to a prompt instruction. Yellow marks user input, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder. The user input and Symbol instance attribute values are evaluated and transformed according to the pre- and post-processing steps. A custom method is constructed by using our reference implementation of decorators and passes the user input and instance attributes to the respective engine. A pre-processor manipulates the string representation of user input and Symbol instance attributes before forwarding them to the engine. For instance, in case the NeSy engine is an LLM (i.e. GPT-4), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine. Other attributes, such as examples or templates, can help pre-define structures for the desired LLM output format. After the result is computed by the engine, it is passed to the post-processors and validated against the constraints.
@@ 6 Evaluation
Figure 5: We evaluate GPT-3.5 Turbo (GPT-3.5), GPT-4 Turbo revision 1106 (GPT-4), Gemini-Pro, LLaMA2-Chat 13B, Mistral 7B and Zephyr 7B on five benchmark categories: 1) Associative Prediction (Association) 2) Multi-modal Binding (Modality) 3) Program Synthesis (Code) 4) Functional Logic Components (Logic) and 5) Hierarchical Computational Graphs (Graphs). We denote the similarity score such that higher is better. The similarity score is measured according to a reference baseline and normalized to exclude noise.
@@ 7.2 Future Work
The goal for Algorithm 1 is to be utilized by an advanced learning agent. This agent, employing reinforcement learning methodologies (Ouyang et al., 2022; Li et al., 2023; Rafailov et al., 2023), could leverage our evaluation measure in existing benchmarks (Milani et al., 2020; Swazinna et al., 2022; Schweighofer et al., 2022) as a means to obtain reward signals to addresses a central problem in reinforcement learning, namely credit assignment (Sutton, 1984; Arjona-Medina et al., 2019; Holzleitner et al., 2020; Patil et al., 2020; Widrich et al., 2021; Dinu et al., 2022). Over time, it aims to develop the ability to autonomously generate its own plans, efficiently schedule tasks and subtasks, and carefully select the most suitable tools for each task. Our protocol lays the groundwork for this agent to learn and expand its base set of capabilities (Amaro et al., 2023), moving towards more sophisticated, self-referential orchestration of multi-step tasks. We’ve already noticed that research is shifting towards this type of methodology (Yuan et al., 2024). Furthermore, in Section 6 we’ve only considered a sequential scheduler. However, our objective is to ultimately assess a non-sequential task execution model, allowing for dynamic insertion and out-of-sequence task execution. In addition, we are interested into exploring similarities of our work with Generative Flow Networks (Bengio et al., 2021a, b; Lahlou et al., 2023).
@@ 7.3 Broader Impact
Furthermore, we believe that many gains in representational stability and consistency may be obtained through multi-modal data training and improved alignment based on operator learning oriented functionalities and workflow-related scoring functionalities, analogous to our introduced quality measure. Gains in representational stability also benefit self-instruction and self-referential sub-process evaluations, which enable the dynamic creation and evaluation of complex hierarchical computational graphs. This will enable online learning models to perform, in real-time, skill acquisition of complex concepts with only one or few examples at inference time. We believe this will enable the creation of autonomously self-evolving cognitive architectures (Langley et al., 2009; Dawid & LeCun, 2023; Sumers et al., 2023). We therefore see an inherent connection to generative design as an analogy for creating coherent and stable ”thought” computational graphs, and believe this paves the path toward broad AI systems (see Section B) and is a requirement for developing General AI agents.
@@ Acknowledgement
We are also grateful to our friends John Chong Min Tan and Tim Scarfe, whose communities have been a hub for exhilarating discussions. Their online presence and engagement have enriched the AI research landscape and broadened our perspectives.
@@ Appendix B Broad AI and Neuro-Symbolic Systems
Figure 6: Hierarchical model of ”cognitive” abilities of AI systems (Chollet, 2019; Hochreiter, 2022). The figure contrasts the emergent paradigm of Broad AI with current Narrow AI systems, showcasing Broad AI’s wider range of capabilities, such as knowledge transfer, interaction, adaptability, robustness, abstraction, advanced reasoning, and efficiency. Broad AI aims to mimic human cognitive adaptability and robustness through advanced methodologies like few-shot learning, self-supervised contrastive learning, and context-sensitive sensory processing. Notably, Broad AI applies principles such as conceptual short-term memory and modern Hopfield networks (Ramsauer et al., 2020) to better integrate context and memory, thus avoiding pitfalls like explaining away and short-cut learning. We acknowledge the potential of NeSy systems as a significant step towards AI systems capable of performing any cognitive task with human-like proficiency.
NeSy methods form the basis for developing new cognitive architectures (Newell & Simon, 1956; Newell et al., 1957; Newell & Simon, 1972; Newell, 1990; Langley et al., 2009; Laird, 2022; Dawid & LeCun, 2023; Sumers et al., 2023; LeCun, 2022; Assran et al., 2023), as illustrated in Figure 7. This hybridization produces computational graphs capable of context-aware learning and reasoning, allowing AI to execute complex tasks with human-like flexibility.